Carlos invested 7000 in a money market mutual fund that pays interest on a daily basis. The balance in his account at the end of 8 mo (245 days) was 7432. Find the effective rate at which Carlos's account earned interest over this period (assume a 365-day year).
Found 2 solutions by addingup, Theo:
Answer by addingup(3677)

A= P(1+r/n)^nt
Amount = Principal(1+rate/periods)^time*periods Since the formula is meant for long periods of time, like years, we don't need this. We are going to substitute in rate/periods= rate/365 and in )^nt we'll use 245
-----------
Now let's substitute with the numbers you were given:
7432= 7000(1+r/365)^245 Subtract 7000 on both sides
432/ (1+r/365)^245 Now take the 245 root
245root(432)= 1+r/365
1.025= 1+r/365 Subtract 1 on both sides
0.025= r/365 We found the annual rate. If you need the DAILY rate, go ahead and divide both sides by 365:
0.025/365= r(daily)
If the annual rate is good enough, then 0.025*100= 2.5% is your annual rate.
Lesson learned here: How do you eat an elephant? One bite at a time.
J

Answer by Theo(13342)

p = present value
f = future value
r = interest rate per time period.
n = number of time periods.

in your problem:

p = 7000
f = 7432
n = 245
r = what you want to find.

the general equation is f = p * (1+r)^n which becomes:

7432 = 7000 * (1+r)^245

divide both sides by 7000 to get:

7432 / 7000 = (1+r)^245

take the 245th root of each side of the equation to get:

(7432/7000)^(1/245) = ((1+r)^245)^(1/245)

this becomes (7432/7000)^(1/245) = 1+r

subtract 1 from both sides of this equation and it becomes:

(7432/7000)^(1/245) - 1 = r

solve for r to get:

r = .0002444578442.

that is the daily rate.

add 1 to this and raise it to the 365th power and then subtract 1 from it and you will get the annual effective rate.

that would be (1.00024458)^365 = 1.093317016 - 1 = .0933170155 * 100% = 9.33170155 percent.

the yearly effective rate is .0933170155.
the yearly effective percentage rate = 9.33170155

any small differences at the tail end of these number is due to calculator rounding of the display.

the yearly nominal rate would be equal to .0002444578442 * 365 = .0892271131.

the annual percentage rate (apr) would be equal to .0892271131 * 100% = 8.92271131 percent.

rounding the annual numbers to something reasonable, you get:

annual effective rate = .0933
annual effective percentage rate = 9.33%

annual nominal rate = .0892
annual nominal percentage rate = 8.92%

these rounded numbers will get you pretty close to what you would get with the unrounded numbers.

for example, the nominal percentage rate is 8.92%.
using that number, i calculated a future value of 7431.86 which is mighty close to 7432.